A bus initially with a speed of 25 m/s slows down with an acceleration of -3 m/s2. How far does it travel before coming to a complete stop?

1
Expert's answer
2022-03-20T18:48:59-0400

The distance traveled by a bus

"d=\frac{v_f^2-v_i^2}{2a}=\frac{0^2-25^2}{2\times (-3)}=104\:\rm m"